# Emergent vorticity asymmetry of one and two-layer shallow water system captured by a next-order balanced model

### v1

@@ 4 Conclusion and outlook
We have shown that SWQG+1 is a PV-based balanced model that can capture many balanced ageostrophic effects of the shallow water model in the finite Rossby number regime. In particular, it can model the evolution of vorticity asymmetry of the shallow water model in its one and multi-layer configurations. We use it to explain the different vorticity skewness in one and two-layer simulations, by showing SWQG+1 can capture the correlation between vorticity, divergence, and height for coherent vortices and baroclinic fronts. It is also well-behaved regarding energy and potential enstrophy, even though we cannot prove exact conservation. This shows that SWQG+1 can be applied to study many GFD problems in the atmosphere and ocean on Earth and other planets. For that, there is much work to be done to adapt and extend the SWQG+1 model. The rest of this section describes some possible applications and the desirable extension to the model.
@@
One central challenge in the parametrization of ocean eddies is to model correctly the mean location and variability of the ocean jets (Chassignet & Marshall, 2008). Thiry et al. (2024) show that SWQG is inaccurate in modeling the jet location of a barotropic wind-driven gyre, perhaps the simplest model of the general ocean circulation. Therefore QG does not capture enough phenomenology to be sufficient for eddy parameterization. However, PV-based thinking and balanced thinking have enabled much progress in eddy parametrization and should not be abandoned. SWQG+1 offers a nice intermediate option. It is still simple like SWQG where one prognostic PV variable governs the entire model evolution. However, it is able to capture the important balanced ageostrophic effects. Future work should aim to replicate the asymmetric lean of the jet in a wind-drive gyre in the SWQG+1 model. Boundary conditions for the SWQG+1 model can be adapted from the boundary condition specifications in Muraki et al. (1999). Furthermore, SWQG+1 needs to be able to deal with vanishing layer height if it were to be applied to study realistic isopycnal ocean models.
@@
The shallow water model, especially the two-layer variant, has been a useful simple model to study the nonlinear and non-smooth effects of moisture in the atmosphere (Lambaerts et al., 2012; Bembenek et al., 2020). SWQG has also been adapted to include moisture effects. In particular, definitions of moist PV have been useful to interpret model results (Lapeyre & Held, 2004; Smith & Stechmann, 2017; Lutsko & Hell, 2021; Brown et al., 2023). SWQG+1, with an appropriate definition of moist PV, can be a useful bridge between these two models. The removal of gravity waves removes an undesirable complication in the study of the large-scale evolution of storm tracks. More broadly, it would be interesting to see if moist effects can be sufficiently captured by inverting from the moist PV. That is, are moist effects fundamentally “balanced” or not?
@@
Coming back down to the ocean bottom. The research on the effect of ocean topography on the coherent structure of the ocean eddy has enjoyed a resurgence (Solodoch et al., 2021; Siegelman et al., 2022; LaCasce et al., 2024). Theoretical understanding of this problem typically comes from SWQG, or the even simpler two-dimensional Euler equation (e.g. Bretherton & Haidvogel, 1976; Siegelman et al., 2022). However, the necessary symmetry between different signed vortices in the SWQG model is an unrealistic deficiency. The study of the realistic formation and destruction of coherent vortices requires numerical simulation using ocean models, which are complex. SWQG+1 and Boussinesq QG+1 can bridge this gap. SWQG+1 fixes the issue of SWQG and allows for the realistic evolution of the coherent vortices. In the meantime, there is still only one variable to predict, PV, and it is possible that many PV-based theoretical thinking can be applied to SWQG+1. Of course, the study of coherent structure on other planets can also benefit from upgrading to the SWQG+1 model (Siegelman & Young, 2023). For these applications, SWQG+1 model needs to be extended and tested on the cases with non-constant $f$ and topography.
@@
[Acknowledgements]We thank Oliver Bühler, Guillaume Lapeyre, and Olivier Pauluis for helpful discussions. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. We thank Shenglong Wang for his work on configuring Dedalus on the NYU Greene cluster.
@@ Appendix A Equivalency of the SWQG+1 potential form to the model of V96
We show that our model is equivalent to the model in Warn et al. (1995) and the extended version with $\beta$ in §2(a) of V96 up to the constant terms in the PV inversions.
First, the two models are the same at the QG level. Then the V96 model inverts for the next-order quantities using their (15-17). We show that the two models are equivalent by recovering (15-17) from the SWQG+1 inversions. V96’s (15) is the next-order PV inversion by noting its left-hand-side is $\mathcal{S}(\Phi^{1})$ while the right-hand-side is equal to the right-hand-side of (20). For (16), we have starting from the left-hand-side
$=\frac{f}{gH}\left\{\textit{Bu}^{-1}\right\}\left[\mathcal{S}(G^{1}_{x})-\mathcal{S}(F^{1}_{y})\right]$ (53)
$=-2J(\Phi^{0}_{x},\Phi^{0}_{y})-\left\{\gamma\right\}\beta\Phi_{y}^{0}$ (54)
where we used the elliptic problem for $F^{1}$ and $G^{1}$. We recovered (16), which is named the Bolin–Charney balance equation (Bolin, 1955; Charney, 1955) and contains the cyclogeostrophic balance. Finally for (17) we have
@@
Of course, this is just the shallow water omega equation.
@@
We first write down the thermal wind balance of the QG-level variables
@@
The “imbalance” equation then states
@@
$=$ $-\left\{\textit{Bu}\right\}H(u^{1}_{2,xx}+v^{1}_{2,xy})-\frac{f}{g^{\prime}}\left(-fu_{2}^{1}-g^{\prime}h^{1}_{2,y}+fu_{1}^{1}\right)$ (63)
which implies

### v3

@@ 4 Conclusion and outlook
We have shown that SWQG+1 is a PV-based balanced model that can capture many balanced ageostrophic effects of the shallow water model in the finite Rossby number regime. In particular, it can model the evolution of vorticity asymmetry of the shallow water model in its one and multi-layer configurations. We use it to explain the different vorticity skewness in one and two-layer simulations, by showing SWQG+1 can capture the correlation between vorticity, divergence, and height for coherent vortices and baroclinic fronts. The energy and potential enstrophy closely track the evolution of the full model, even though we cannot prove exact conservation. This shows that SWQG+1 has potential as a tool for the study of many GFD problems in the atmosphere and ocean on Earth and other planets. For that, there is much work to be done to adapt and extend the SWQG+1 model. The rest of this section describes some possible applications and the desirable extensions to the model.
@@
One central challenge in the parametrization of ocean eddies is to model correctly the mean location and variability of the ocean jets (Chassignet & Marshall, 2008). Thiry et al. (2024) show that SWQG is inaccurate in modeling the jet location of a barotropic wind-driven gyre, perhaps the simplest model of the general ocean circulation. Therefore, QG does not capture enough phenomenology to be sufficient for eddy parameterization. However, PV-based thinking and balanced thinking have enabled much progress in eddy parametrization and should not be abandoned. SWQG+1 offers a nice intermediate option. It is still simple like SWQG where one prognostic PV variable governs the entire model evolution. However, it is able to capture the important balanced ageostrophic effects. Future work should aim to replicate the asymmetric lean of the jet in a wind-driven gyre in the SWQG+1 model. Boundary conditions for the SWQG+1 model can be adapted from the boundary condition specifications in Muraki et al. (1999).
@@
The shallow water model in one-layer (Lahaye & Zeitlin, 2016) and two-layer (Lambaerts et al., 2012; Bembenek et al., 2020) has been a useful simple model to study the nonlinear and non-smooth effects of moisture in the mid-latitude atmosphere. SWQG has also been adapted to include moisture effects. In particular, definitions of moist PV have been useful to interpret model results (Lapeyre & Held, 2004; Smith & Stechmann, 2017; Lutsko & Hell, 2021; Brown et al., 2023). SWQG+1, with an appropriate definition of moist PV, can be a useful bridge between these two models. The removal of gravity waves removes an undesirable complication in the study of the large-scale evolution of storm tracks. More broadly, it would be interesting to see if moist effects can be sufficiently captured by inverting from the moist PV. That is, are moist effects fundamentally “balanced” or not?
@@
The research on the effect of ocean topography on the coherent structure of the ocean eddy has enjoyed a resurgence (Solodoch et al., 2021; Siegelman & Young, 2023; LaCasce et al., 2024). Theoretical understanding of this problem typically comes from SWQG, or the even simpler two-dimensional Euler equation (e.g. Bretherton & Haidvogel, 1976; Siegelman & Young, 2023). However, the necessary symmetry between different signed vortices in the SWQG model is unrealistic. The study of the realistic formation and destruction of coherent vortices requires numerical simulation using complex ocean models. SWQG+1 and Boussinesq QG+1 can bridge this gap. SWQG+1 improves upon SWQG and allows for the modeling of realistic evolution of the coherent vortices (though it still assumes asymptotically small topography height). In the meantime, there is still only one variable to predict, PV, and it is possible that many PV-based theoretical thinking can be applied to SWQG+1. Of course, the study of coherent structure on other planets can also benefit from upgrading to the SWQG+1 model (Siegelman et al., 2022). For these applications, SWQG+1 model needs to be extended and tested on the cases with non-constant $f$ and topography.
@@
[Acknowledgements]We thank Oliver Bühler, Guillaume Lapeyre, and Olivier Pauluis for helpful discussions. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. We thank Shenglong Wang for his work on configuring Dedalus on the NYU Greene cluster. We thank three anonymous reviewers for their helpful suggestions.
@@ Appendix A Equivalency of the SWQG+1 potential form to the model of V96
We show that our model is equivalent to the model in Warn et al. (1995) and the extended version under the $\beta$-approximation in §2(a) of V96 up to the constant terms in the PV inversions.
First, the two models are the same at the QG level. Then the V96 model inverts for the next-order quantities using their (V15-V17). We show that the two models are equivalent by recovering (V15-V17) from the SWQG+1 inversions. V96’s (V15) is the next-order PV inversion by noting its left-hand-side is $\mathcal{S}(\Phi^{1})$ while the right-hand-side is equal to the right-hand-side of (20). For (V17), the shallow water $\omega$-equation, we have
@@
To recover (V16), the Bolin–Charney balance equation (Bolin, 1955; Charney, 1955), we have
=$=\left\{\textit{Bu}\right\}\frac{gH}{f}\left[\mathcal{S}(G^{1}_{x})-\mathcal{S}(F^{1}_{y})\right]$ (57)
$=-2J(\Phi^{0}_{x},\Phi^{0}_{y})-\left\{\gamma\right\}\beta\Phi_{y}^{0}-\left\{\gamma\right\}\beta y\nabla^{2}\Phi^{0}$ (58)
where we used the elliptic problem for $F^{1}$ and $G^{1}$. The last term on the right-hand side is not included in (V16). This is a mistake in V96 which stems from the same term missing in the “divergence equation” (V6). Forming the divergence of the shallow water momentum equations (1) reveals (V6) should additionally include a $-\beta y\zeta$ term.
@@
We start with the two-layer shallow water system (36) and the potential form for the physical variables in the two-layer case (41). For the QG-level variables, they are in thermal wind balance. For $h_{1,x}$, this is
@@
The “imbalance” equation is the time evolution equation of the difference between the thermally balanced variables, that is, the “imbalance”. For the QG-level geostrophically balanced variables, the difference should be zero for all time. This allows us to form a diagnostic equation free of time partial derivatives.
@@
$=$ $-\left\{\textit{Bu}\right\}H(u^{1}_{2,xx}+v^{1}_{2,xy})-\frac{f}{g^{\prime}}\left(-fu_{2}^{1}-g^{\prime}h^{1}_{2,y}+fu_{1}^{1}\right).$ (65)
Similar steps gives